OpenStudy (anonymous):

find the domain of the function g(x)=1/10-9x Can someone please help

OpenStudy (anonymous):

seeing 0 cant be in the denominator we have to exclude 0 from the domain so, g(x) = 1/10-9x = 10-9x = 0 => 10 = 9x => 10/9 = x so with this given value in x its determined that the denominator will be 0 at 10/9 so to exclude this from the domain we write the domain like this \[(-\infty,10/9)U(10/9,\infty)\] what this is saying is that the domain consists of any number between - infinity to positive infinity excluding 10/9

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you're quite welcome

OpenStudy (anonymous):

the domain is R/{10/9}